Find the unit vector parellal to the vector A=2i+j+k and B=i-k

Grade:11

3 Answers

Nishant Vora IIT Patna
8 years ago
A cap =(2i+j+k) / root6 and B=(i-k)/ root2

Thanks
SREEKANTH INDLA
8 years ago
hai friends see the answer
unit vocter perpendicular A and B
A*B=-I+3J-K
A*B=ROOT 11
n=a*b/a*b
-i+3j-k/root 11
K RAMBABU
8 years ago
according to poblem 
unit vector perpendicular A and B
A*B=-i+3j-k
A*B=root of 11
n=A*B/A*B
-I+3J-K/ROOT OF11